# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ Acknowledgements.
S2.2.1 Quantum Approximate Optimization Algorithm
Description
@@
$\hat{\tilde{C}}=\frac{1}{2}\sum\nolimits_{i,j=1}^{N}\mathsf{W}_{ij}\hat{Z}_{i}\hat{Z}_{j},$ (S12)
@@
$\bigl{|}\Psi\bigr{\rangle}_{p}=\left[\prod\nolimits_{\ell=1}^{p}e^{-i\beta_{\ell}\sum_{j=1}^{N}\hat{X}_{j}}e^{i\gamma_{\ell}\hat{\tilde{C}}}\right]\hat{H}^{\otimes N}|{0}\rangle^{\otimes N},$ (S13)
where $\hat{X}_{i}$ is the Pauli operator on qubit $i$, $\hat{H}$ is the one-qubit Hadamard gate, the operator $\hat{\tilde{C}}$ is defined in Eq. (S12), and $\{\gamma_{\ell},\beta_{\ell}\}$ are real-valued angles. Ideally, these angles should be chosen such that they extremize the expectation value $\langle\hat{\tilde{C}}\rangle_{p}$ over the quantum state $|\Psi\rangle_{p}$ of Eq. (S13). The first term of the bracket is known as the mixer, and the second one as the phase separator.
Remarks
@@
Here, we focus on unit-weight random $3$-regular graph problem instances. We rely on the fixed-angle conjecture of Refs. [35, 36] to set the QAOA angles $\{\boldsymbol{\gamma},\boldsymbol{\beta}\}_{p}$, thus strictly avoiding any optimization of the QAOA angles while maintaining near-optimality. We use the values reported in Ref. [36], and reproduced in Tab. S2.
@@
Table S2: QAOA angles for unit-weight $3$-regular graphs up to $p=3$ QAOA layers, as defined in Eq. (S13), extracted from Ref. [36]. We use these angles throughout this work, unless specified otherwise. Units are radians.
Algorithmic Complexity
The execution time of the QAOA to obtain $|\Psi\rangle_{p}$ from Eq. (S13) is linear with the number of layers $p$. Sampling $n_{\textrm{ex}}$ bit strings from the resulting quantum state $|\Psi\rangle_{p}$ is also linear in $n_{\textrm{ex}}$. The algorithmic complexity scaling of the QAOA with the number of qubits $N$ depends on the implementation: quantum simulations versus classical emulations, and the nature of the emulator and quantum hardware, etc. We discuss the specific cases of interest in Sec. S4.
S2.2.2 Quantum-Flavored Relax-and-Round Algorithm
Description
@@
$\mathsf{Z}_{ij}^{(p)}=\bigl{(}\delta_{ij}-1\bigr{)}\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}_{p},$ (S14)
where $\hat{Z}_{i}$ is the Pauli operator on qubit $i$, $\delta_{ij}$ is the Kronecker delta, and the expectation value $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ is taken over the quantum state $|\Psi\rangle_{p}$ resulting from the QAOA of Eq. (S13). $\mathsf{Z}^{(p)}$ is an $N\times N$ real and symmetric matrix with null entries on its diagonal. The next step consists of running the classical relax-and-round step explained in Sec. S2.1.7 on the correlation matrix $\mathsf{Z}^{(p)}$ instead of the matrix $\mathsf{W}$ encoding the problem.
Remarks
@@
$\mathsf{Z}^{(p=\infty)}=\mathsf{I}-\boldsymbol{z}_{\textrm{opt}}\otimes\boldsymbol{z}_{\textrm{opt}},$ (S15)
where $\otimes$ denotes the outer product. The first term of Eq. (S15) is the identity matrix. The second term is a rank one matrix with eigenvector $\propto\boldsymbol{z}_{\textrm{opt}}$. A classical relax-and-round step on the correlation matrix of Eq. (S15) will therefore return the optimal solution $\boldsymbol{z}_{\textrm{opt}}$. Ref. [27] provides a strategy for dealing with degenerate optimal solutions.
@@
Algorithmic Complexity
Compared to the classical relax-and-round algorithm of Sec. S2.1.7, the algorithmic complexity of the quantum version has the additional cost of obtaining the correlation matrix $\mathsf{Z}^{(p)}$.
Moreover, the density of the correlation matrix is going to be larger than that of the adjacency matrix, $\mathsf{W}$. While in the limit of large $N$, one still expects $O(1)$ nonzero entries per row (see Sec. S3.2), at intermediate $N$, and depending on the number of layers $p$, the correlation matrix will get denser. As such, the scaling of the eigendecomposition will tend towards $O(N^{2})$ instead of $O(N)$.
S2.2.3 Quantum-Flavored Goemans-Williamson Algorithm
Description
Ref. [27] introduced a quantum version of the classical Goemans-Williamson algorithm (QGW) outlined in Sec. S2.1.8. The first step is to construct the correlation matrix $\mathsf{Z}^{(p)}$ of Eq. (S14) using the QAOA with $p$ layers. Next, one substitutes the matrix $\mathsf{W}$ encoding the problem in Eq. (S11) with the correlation matrix $\mathsf{Z}^{(p)}$. The eigenvalue-based Goemans-Williamson algorithm of Sec. S2.1.8 is then executed as previously described.
Remarks
The QGW algorithm yields the optimal solution $\boldsymbol{z}_{\textrm{opt}}$ in the $p=\infty$ limit  [27]. This can be shown by substituting the $p\to+\infty$ correlation matrix of Eq. (S15) into Eq. (S11) and choosing $\mathsf{diag}(\boldsymbol{u})=\textrm{tr}(\mathsf{D})\mathsf{I}/N-\mathsf{D}$, thereby ensuring the leading eigenvector is $\propto\boldsymbol{z}_{\textrm{opt}}$. In  [27] it was shown analytically that at $p=1$ the performance of the QGW algorithm was on par with that of the classical Goemans-Williamson algorithm for ring and complete graphs. Furthermore, numerical tests at $p=1$ on several instances of other non-trivial problems, including unit-weight random $3$-regular graphs, showed the quantum and classical algorithms to perform similarly. It should be noted that the average performance of the QGW algorithm increases with the number of layers $p$, until convergence to the optimal solution for $p\to+\infty$ layers.
Algorithmic Complexity
Compared to the classical Goemans-Williamson algorithm of Sec. S2.1.8, the algorithmic complexity of the quantum version has the additional cost of obtaining the correlation matrix $\mathsf{Z}^{(p)}$.
S2.2.4 Greedy-Enhanced Relax-and-Round Algorithm
Description
@@
$p_{i}=|{z_{i}}|^{-1}\Bigr{/}\sum\nolimits_{i=1}^{N}|{z_{i}}|^{-1},$ (S16)
such that variables $z_{i}$ closer to zero have a larger chance of being visited. The choice of mapping between $p_{i}$ and $z_{i}$ is nonunique and could probably be fine-tuned. We set the number of variables visited as $fN$ with $f=10$ for an $N$-variable problem. We study the performance of this greedy heuristic as a function of $f$ in Sec. S5.7.
Remarks
We note again that this strategy is not specific to quantum-flavored relax-and-round solvers but could apply to any. Such a step is usually included in solvers, given its low algorithmic overhead. We investigate different sign-flip strategies in Sec. S5.7. In addition, we study the performance gain of the greedy-enhanced quantum relax-and-round solver over its original version in Sec. S5.7.
Algorithmic Complexity
@@
S3 Problem Instances
S3.1 Unit-weight Random $3$-Regular Graphs
Figure S2: Average shortest path length in unit-weight $3$-regular random graphs as a function of the number of vertices $N$. Average taken over all pairs of vertices in a graph and over $50$ to $1,000$ randomly generated graphs. The solid line is a fit of the form $A\ln(N/B)$, where $A$ and $B$ are real-valued parameters obtained by least-square fitting. Error bars indicate one standard deviation.
In this work, we focus on $N$-variable problem instances defined by Eq. (S1). Each instance is characterized by an $N\times N$ matrix $\mathsf{W}$, drawn uniformly from the ensemble of adjacency matrices of unit-weight $3$-regular graphs [80, 81]. $3$-regular graphs are such that their vertices are connected to three other unique vertices. Therefore, $\mathsf{W}$ is sparse, real, symmetric, and contains only three nonzero entries per row and column, all equal to unity, for a total of $3N$ entries. By definition, $N\geq 4$ and $N$ is even. The size of the ensemble at fixed $N$ is given by Ref. [82, A008277]. We generate the problem instances using the Python package NetworkX [83]. In the limit of large $N$, the average shortest path between two vertices grows as $\ln N$, see Fig. S2. Moreover, the statistics on the cycles in a typical $3$-regular graph are known. The probability to have $K$ cycles of length $\ell$ is Poissonian and given by [84, 85, 38]
$p(K,\ell)=\frac{\lambda^{K}e^{-\lambda}}{K!}\quad\textrm{where}~{}\lambda=\frac{2^{\ell}}{2\ell}.$ (S17)
@@
$\mathbb{E}\bigl{[}K_{\ell}\bigr{]}=\sum\nolimits_{K=1}^{\infty}Kp(K,\ell)=\lambda=\frac{2^{\ell}}{2\ell},$ (S18)
@@
S3.2 Correlations are Constrained by the QAOA Light Cone
S3.2.1 The Light Cone Picture
Figure S3: (a-d) An $N=10$ variable $3$-regular graph. An arbitrary reference qubit is colored in teal. Its $p$-nearest neighbors are colored in magenta. (a) $p=1$, (b) $p=2$, and (c) $p=3$. (d) Two arbitrary reference qubits with their nearest neighbors ($p=1$). These two qubits have intersecting light cones. (e) Qubits involved in two intersecting light cones made of two $3$-Cayley trees connected through a single common qubit.
The quantum algorithms that we leverage are based on the expectation value of two-point correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ between all problem variables $i$ and $j$ to build the correlation matrix $\mathsf{Z}^{(p)}$. Naively, this implies computing $N(N-1)/2\sim O(N^{2})$ elements. A QAOA circuit with $p\geq 1$ layers, as defined in Eq. (S13), generates a light cone for each of the qubits centered around that qubit and including all its $p$-nearest neighbors, see Figs. S3a, S3b, and S3c. Thus, for a $3$-regular graph, such a light cone will include at most
$1+3(2^{p}-1),$ (S19)
@@
The existence of a light cone is such that the correlation $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ is trivially zero if the light cones centered around qubits $i$ and $j$, do not intersect. In the limit of large $N$, because the average shortest path between two vertices grows as $\ln N$ (see Fig. S2), this means that for a number of QAOA layers $p\ll\ln N$, one expects only $O(N)$ elements to be nonzero. For a given $3$-regular graph problem instance, these $O(N)$ pairs $(i,j)$ with a nonzero correlation can be found in $O(pN^{2})$ operations. We show an example in Fig. S3d of two intersecting light cones at $p=1$.
S3.2.2 Implications for the QAOA Runs
@@
Subgraph size distribution
Figure S4: Distribution of the size of the subgraphs from intersecting light cones, for different problem sizes $N$. First row: $p=1$ QAOA layer. Second row: $p=2$ QAOA layers. Third row: $p=3$ QAOA layers. Columns from left to right correspond to: $N=64$, $N=256$, $N=1024$, $N=2048$, and $N=4096$. The maximum size is independent of the problem size $N$ at large $N$ and bounded by Eq. (S20).
In Fig. S4, we show for different problem sizes $N$ the distribution of the size of the subgraphs generated from the intersection of the light cones on two variables. The maximum size depends on $p$ and is bounded by Eq. (S20). At large $N$, the distribution suggests that large subgraphs are predominant.
Subgraph treewidth distribution
Figure S5: Distribution of the treewidth of the subgraphs from intersecting light cones, for different problem sizes $N$. First row: $p=1$ QAOA layer. Second row: $p=2$ QAOA layers. Third row: $p=3$ QAOA layers. Columns from left to right correspond to: $N=64$, $N=256$, $N=1024$, $N=2048$, and $N=4096$.
In addition to the size of the subgraphs, we consider their treewidth, which we compute using the minimum fill-in heuristic [86, 83]. A treewidth of one means that the graph is a tree. As the treewidth increases, the graph becomes less tree-like and includes loops. We show in Fig. S5 the subgraph treewidth distribution for different number of QAOA layers $p$ and problem sizes $N$. As $N$ increases, we observe that the distribution narrows and peaks at a treewidth of one. This suggests that the majority of the subgraphs are trees.
Fig. S4 shows that the majority of the subgraphs saturate the maximum size bound of Eq. (S20). We, therefore, conclude that most of the subgraphs consist of two $3$-Cayley trees connected by a single common vertex.
S3.3 Graph Isomorphism
S3.3.1 Connection to the Light Cone Picture from the QAOA
When using the light cone approach (see Sec. S3.2), a single QAOA run on the complete graph is replaced by $O(N)$ QAOA runs on subgraphs formed by the intersecting light cones. Of the $O(N)$ subgraphs evaluated, we posit that the number of unique subgraphs is independent of $N$ in the large $N$ limit. Said differently, there are $O(N)$ isomorphic graphs formed by the intersecting light cones. Hence, determining the isomorphism can reduce to $O(1)$ the number of QAOA runs required to compute $\mathsf{Z}$ for a number of QAOA layers $p\ll\ln N$ in the limit of large $N$. This assertion is based on the tendency for subgraphs to become more tree-like with increasing $N$ (see the treewidth analysis in Sec. S3.2.2) and supported by the numerical data presented in the remainder of this Section.
Working with an Isomorphism Database

### v2

@@ B.2 Quantum
B.2.1 Quantum Approximate Optimization Algorithm
Description—
@@ B.2.1 Quantum Approximate Optimization Algorithm
$\hat{\tilde{C}}=\frac{1}{2}\sum\nolimits_{i,j=1}^{N}\mathsf{W}_{ij}\hat{Z}_{i}\hat{Z}_{j},$ (15)
@@
$\bigl{|}\Psi\bigr{\rangle}_{p}=\left[\prod\nolimits_{\ell=1}^{p}e^{-i\beta_{\ell}\sum_{j=1}^{N}\hat{X}_{j}}e^{i\gamma_{\ell}\hat{\tilde{C}}}\right]\hat{H}^{\otimes N}|{0}\rangle^{\otimes N},$ (16)
where $\hat{X}_{i}$ is the Pauli operator on qubit $i$, $\hat{H}$ is the one-qubit Hadamard gate, the operator $\hat{\tilde{C}}$ is defined in Eq. (15), and $\{\gamma_{\ell},\beta_{\ell}\}$ are real-valued angles. Ideally, these angles should be chosen such that they extremize the expectation value $\langle\hat{\tilde{C}}\rangle_{p}$ over the quantum state $|\Psi\rangle_{p}$ of Eq. (16). The first term of the bracket is known as the mixer, and the second one as the phase separator.
Remarks—
@@
Here, we focus on unit-weight random $3$-regular graph problem instances. We rely on the fixed-angle conjecture of Refs. [31, 25] to set the QAOA angles $\{\boldsymbol{\gamma},\boldsymbol{\beta}\}_{p}$, thus strictly avoiding any optimization of the QAOA angles while maintaining near-optimality. We use the values reported in Ref. [25], and reproduced in Tab. 2.
@@
Table 2: QAOA angles for unit-weight $3$-regular graphs up to $p=3$ QAOA layers, as defined in Eq. (16), extracted from Ref. [25]. We use these angles throughout this work, unless specified otherwise. Units are radians.
Algorithmic Complexity—
The execution time of the QAOA to obtain $|\Psi\rangle_{p}$ from Eq. (16) is linear with the number of layers $p$. Sampling $n_{\textrm{ex}}$ bit strings from the resulting quantum state $|\Psi\rangle_{p}$ is also linear in $n_{\textrm{ex}}$. The algorithmic complexity scaling of the QAOA with the number of qubits $N$ depends on the implementation: quantum simulations versus classical emulations, and the nature of the emulator and quantum hardware, etc. We discuss the specific cases of interest in App. D.
B.2.2 Quantum-Flavored Relax-and-Round Algorithm
Description—
@@ B.2.2 Quantum-Flavored Relax-and-Round Algorithm
$\mathsf{Z}_{ij}^{(p)}=\bigl{(}\delta_{ij}-1\bigr{)}\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}_{p},$ (17)
where $\hat{Z}_{i}$ is the Pauli operator on qubit $i$, $\delta_{ij}$ is the Kronecker delta, and the expectation value $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ is taken over the quantum state $|\Psi\rangle_{p}$ resulting from the QAOA of Eq. (16). $\mathsf{Z}^{(p)}$ is an $N\times N$ real and symmetric matrix with null entries on its diagonal. The next step consists of running the classical relax-and-round step explained in App. B.1.7 on the correlation matrix $\mathsf{Z}^{(p)}$ instead of the matrix $\mathsf{W}$ encoding the problem.
Remarks—
@@
$\mathsf{Z}^{(p=\infty)}=\mathsf{I}-\boldsymbol{z}_{\textrm{opt}}\otimes\boldsymbol{z}_{\textrm{opt}},$ (18)
where $\otimes$ denotes the outer product. The first term of Eq. (18) is the identity matrix. The second term is a rank one matrix with eigenvector $\propto\boldsymbol{z}_{\textrm{opt}}$. A classical relax-and-round step on the correlation matrix of Eq. (18) will therefore return the optimal solution $\boldsymbol{z}_{\textrm{opt}}$. Ref. [29] provides a strategy for dealing with degenerate optimal solutions.
@@
Algorithmic Complexity—
Compared to the classical relax-and-round algorithm of App. B.1.7, the algorithmic complexity of the quantum version has the additional cost of obtaining the correlation matrix $\mathsf{Z}^{(p)}$.
Moreover, the density of the correlation matrix is going to be larger than that of the adjacency matrix, $\mathsf{W}$. While in the limit of large $N$, one still expects $O(1)$ nonzero entries per row (see App. C.2), at intermediate $N$, and depending on the number of layers $p$, the correlation matrix will get denser. As such, the scaling of the eigendecomposition will tend toward $O(N^{2})$ instead of $O(N)$.
B.2.3 Quantum-Flavored Goemans-Williamson Algorithm
Description—
Ref. [29] introduced a quantum version of the classical Goemans-Williamson algorithm (QGW) outlined in App. B.1.8. The first step is to construct the correlation matrix $\mathsf{Z}^{(p)}$ of Eq. (17) using the QAOA with $p$ layers. Next, one substitutes the matrix $\mathsf{W}$ encoding the problem in Eq. (14) with the correlation matrix $\mathsf{Z}^{(p)}$. The eigenvalue-based Goemans-Williamson algorithm of App. B.1.8 is then executed as previously described.
Remarks—
The QGW algorithm yields the optimal solution $\boldsymbol{z}_{\textrm{opt}}$ in the $p=\infty$ limit  [29]. This can be shown by substituting the $p\to+\infty$ correlation matrix of Eq. (18) into Eq. (14) and choosing $\mathsf{diag}(\boldsymbol{u})=\textrm{tr}(\mathsf{D})\mathsf{I}/N-\mathsf{D}$, thereby ensuring the leading eigenvector is $\propto\boldsymbol{z}_{\textrm{opt}}$. In  [29] it was shown analytically that at $p=1$ the performance of the QGW algorithm was on par with that of the classical Goemans-Williamson algorithm for ring and complete graphs. Furthermore, numerical tests at $p=1$ on several instances of other nontrivial problems, including unit-weight random $3$-regular graphs, showed the quantum and classical algorithms to perform similarly. It should be noted that the average performance of the QGW algorithm increases with the number of layers $p$, until convergence to the optimal solution for $p\to+\infty$ layers.
Algorithmic Complexity—
Compared to the classical Goemans-Williamson algorithm of App. B.1.8, the algorithmic complexity of the quantum version has the additional cost of obtaining the correlation matrix $\mathsf{Z}^{(p)}$.
B.2.4 Greedy-Enhanced Relax-and-Round Algorithm
Description—
@@ B.2.4 Greedy-Enhanced Relax-and-Round Algorithm
$p_{i}=|{z_{i}}|^{-1}\Bigr{/}\sum\nolimits_{i=1}^{N}|{z_{i}}|^{-1},$ (19)
such that variables $z_{i}$ closer to zero have a larger chance of being visited. The choice of mapping between $p_{i}$ and $z_{i}$ is nonunique and could probably be fine-tuned. We set the number of variables visited as $fN$ with $f=10$ for an $N$-variable problem. We study the performance of this greedy heuristic as a function of $f$ in App. E.6.
Remarks—
We note again that this strategy is not specific to quantum-flavored relax-and-round solvers but could apply to any. Such a step is usually included in solvers, given its low algorithmic overhead. We investigate different sign-flip strategies in App. E.6. In addition, we study the performance gain of the greedy-enhanced quantum relax-and-round solver over its original version in App. E.6.
Algorithmic Complexity—
@@
Appendix C Problem Instances
C.1 Unit-weight Random $3$-Regular Graphs
Figure 5: Average shortest path length in unit-weight $3$-regular random graphs as a function of the number of vertices $N$. Average taken over all pairs of vertices in a graph and over $50$ to $1,000$ randomly generated graphs. The solid line is a fit of the form $A\ln(N/B)$, where $A$ and $B$ are real-valued parameters obtained by least-square fitting. Error bars indicate one standard deviation.
In this work, we focus on $N$-variable problem instances defined by Eq. (4). Each instance is characterized by an $N\times N$ matrix $\mathsf{W}$, drawn uniformly from the ensemble of adjacency matrices of unit-weight $3$-regular graphs [78, 79]. $3$-regular graphs are such that their vertices are connected to three other unique vertices. Therefore, $\mathsf{W}$ is sparse, real, symmetric, and contains only three nonzero entries per row and column, all equal to unity, for a total of $3N$ entries. By definition, $N\geq 4$ and $N$ is even. The size of the ensemble at fixed $N$ is given by Ref. [80, A008277]. We generate the problem instances using the Python package NetworkX [81]. In the limit of large $N$, the average shortest path between two vertices grows as $\ln N$, see Fig. 5. Moreover, the statistics on the cycles in a typical $3$-regular graph are known. The probability to have $K$ cycles of length $\ell$ is Poissonian and given by [82, 83, 38]
$p(K,\ell)=\frac{\lambda^{K}e^{-\lambda}}{K!}\quad\textrm{where}~{}\lambda=\frac{2^{\ell}}{2\ell}.$ (20)
@@ C.1 Unit-weight Random $3$-Regular Graphs
$\mathbb{E}\bigl{[}K_{\ell}\bigr{]}=\sum\nolimits_{K=1}^{\infty}Kp(K,\ell)=\lambda=\frac{2^{\ell}}{2\ell},$ (21)
@@
C.2 Correlations are Constrained by the QAOA Light Cone
C.2.1 The Light Cone Picture
Figure 6: (a-d) An $N=10$ variable $3$-regular graph. An arbitrary reference qubit is colored in teal. Its $p$-nearest neighbors are colored in magenta. (a) $p=1$, (b) $p=2$, and (c) $p=3$. (d) Two arbitrary reference qubits with their nearest neighbors ($p=1$). These two qubits have intersecting light cones. (e) Qubits involved in two intersecting light cones made of two $3$-Cayley trees connected through a single common qubit.
The quantum algorithms that we leverage are based on the expectation value of two-point correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ between all problem variables $i$ and $j$ to build the correlation matrix $\mathsf{Z}^{(p)}$. Naively, this implies computing $N(N-1)/2\sim O(N^{2})$ elements. A QAOA circuit with $p\geq 1$ layers, as defined in Eq. (16), generates a light cone for each of the qubits centered around that qubit and including all its $p$-nearest neighbors, see Figs. 6a, 6b, and 6c. Thus, for a $3$-regular graph, such a light cone will include at most
$1+3(2^{p}-1),$ (22)
@@ C.2.1 The Light Cone Picture
The existence of a light cone is such that the correlation $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ is trivially zero if the light cones centered around qubits $i$ and $j$, do not intersect. In the limit of large $N$, because the average shortest path between two vertices grows as $\ln N$ (see Fig. 5), this means that for a number of QAOA layers $p\ll\ln N$, one expects only $O(N)$ elements to be nonzero. For a given $3$-regular graph problem instance, these $O(N)$ pairs $(i,j)$ with a nonzero correlation can be found in $O(pN^{2})$ operations. We show an example in Fig. 6d of two intersecting light cones at $p=1$.
C.2.2 Implications for the QAOA Runs
@@
Subgraph size distribution—
Figure 7: Distribution of the size of the subgraphs from intersecting light cones, for different problem sizes $N$. First row: $p=1$ QAOA layer. Second row: $p=2$ QAOA layers. Third row: $p=3$ QAOA layers. Columns from left to right correspond to: $N=64$, $N=256$, $N=1024$, $N=2048$, and $N=4096$. The maximum size is independent of the problem size $N$ at large $N$ and bounded by Eq. (23).
In Fig. 7, we show for different problem sizes $N$ the distribution of the size of the subgraphs generated from the intersection of the light cones on two variables. The maximum size depends on $p$ and is bounded by Eq. (23). At large $N$, the distribution suggests that large subgraphs are predominant.
Subgraph treewidth distribution—
Figure 8: Distribution of the treewidth of the subgraphs from intersecting light cones, for different problem sizes $N$. First row: $p=1$ QAOA layer. Second row: $p=2$ QAOA layers. Third row: $p=3$ QAOA layers. Columns from left to right correspond to: $N=64$, $N=256$, $N=1024$, $N=2048$, and $N=4096$.
In addition to the size of the subgraphs, we consider their treewidth, which we compute using the minimum fill-in heuristic [84, 81]. A treewidth of one means that the graph is a tree. As the treewidth increases, the graph becomes less tree-like and includes loops. We show in Fig. 8 the subgraph treewidth distribution for different number of QAOA layers $p$ and problem sizes $N$. As $N$ increases, we observe that the distribution narrows and peaks at a treewidth of one. This suggests that the majority of the subgraphs are trees.
Fig. 7 shows that the majority of the subgraphs saturate the maximum size bound of Eq. (23). We, therefore, conclude that most of the subgraphs consist of two $3$-Cayley trees connected by a single common vertex.
C.3 Graph Isomorphism
C.3.1 Connection to the Light Cone Picture from the QAOA
When using the light cone approach (see App. C.2), a single QAOA run on the complete graph is replaced by $O(N)$ QAOA runs on subgraphs formed by the intersecting light cones. Of the $O(N)$ subgraphs evaluated, we posit that the number of unique subgraphs is independent of $N$ in the large $N$ limit. Said differently, there are $O(N)$ isomorphic graphs formed by the intersecting light cones. Hence, determining the isomorphism can reduce to $O(1)$ the number of QAOA runs required to compute $\mathsf{Z}$ for a number of QAOA layers $p\ll\ln N$ in the limit of large $N$. This assertion is based on the tendency for subgraphs to become more tree-like with increasing $N$ (see the treewidth analysis in App. C.2.2) and supported by the numerical data presented in the remainder of this Section.
Working with an Isomorphism Database—